Solve the following. eliminate extraneous solutions.
a 7^2x+1=4^x-2
b. 4^x= 3+18(4^-x)
c. 9= log _5(x+100)+6

Answer by Alan3354(69443)

a 7^2x+1=4^x-2
Ambiguous.
Is the exponent of 7 2x? or 2x+1?
-------------
b. 4^x= 3+18(4^-x)
No algebraic solution.
Use graphical or iterative methods.
---------------
c. 9= log _5(x+100)+6
What does the underscore mean?